OpenStudy (anonymous):

Find the domain \[f(x)=\sqrt{2x-1}\]

OpenStudy (anonymous):

solve\[2x+1 \ge 0\]

OpenStudy (anonymous):

the number inside the radical must be greater than or equal to 0. So you do this \[2x - 1 \ge 0\] Solve for x and you get \[x \ge \frac{1}{2}\] Therefore the domain is \[[\frac{1}{2},\infty)\]

OpenStudy (anonymous):

and how about just a simple \[f(x)=x ^{3}-2\] ?

OpenStudy (anonymous):

^^ all real numbers

OpenStudy (anonymous):

so confused... but thank you!

OpenStudy (anonymous):

What are you confused about?

OpenStudy (anonymous):

just what im supposed to do, i have a test tomorrow. theres one such as f(x)=\[1divx ^{2}+5x+6\]

OpenStudy (ybarrap):

just look for real numbers that are "legal' under the radical. In this case, you can't have negative square roots.

OpenStudy (anonymous):

this server is killing me

OpenStudy (anonymous):

\[\frac{1}{x^2 + 5x + 6}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok you will need to find what number(s) make the denominator 0 because division by 0 is undefined. So set \[x^2 + 5x + 6 = 0\] and solve for x

OpenStudy (ybarrap):

For fractions, you can't have zeros in the denominator. What are the roots of x^2 + 5x + 6? These are where the domain make the fraction undefined (i.e. 1/0)

OpenStudy (anonymous):

thank youu!

OpenStudy (ybarrap):

(x+3)(x+2) so zero occurs at x=-3 and x=-2. So domain is all real except -3 and -2.

OpenStudy (anonymous):

and for square roots, make it negative?

OpenStudy (ybarrap):

yes, no negatives allowed in square roots. That's it. That's all the rules for determining domains. :)

OpenStudy (anonymous):

for square roots the number inside the radical MUST be greater or equal to 0.. Because if you take a square root of a negative number you get imaginary/complex numbers

OpenStudy (anonymous):

thanks =)

OpenStudy (anonymous):

So all you do is set what ever is inside the radical greater or equal to 0 and solve for the variable.

OpenStudy (anonymous):

Thats all the rules.. No division by 0 and no negative square roots.

OpenStudy (ybarrap):

We want real solutions. In Complex domains, negative square roots are allowed and division by zero is also allowed (only if imaginary part is not) -- but that's an entirely other great story.

OpenStudy (anonymous):

^ thats a ugly story lol :P

OpenStudy (anonymous):

mhmk ill figure it out, thanks for all the help guys

OpenStudy (ybarrap):

np - keep it up!